# Robust Universal Inference For Misspecified Models

### v3

@@ 10 Discussion
The papers [58, 23] study many variants of universal inference (including constructing confidence sequences instead of confidence sets, to combining multiple sample-splits) and investigating these variants in the context of the robust universal inference framework of this paper would be interesting.
@@ Appendix A Examples of IPM-ReDI set
Suppose $\rho(P_{0}\|P_{1})=\mathsf{TV}(P_{0},P_{1})$ where $\mathsf{TV}$ is the total variation distance. This is an IPM over the function class =$\mathcal{F}=\{f:\left\|f\right\|_{\infty}\leq 1\}$. An equivalent definition is $\mathsf{TV}(P_{0},P_{1})=\sup_{A}|P_{0}(A)-P_{1}(A)|=P_{0}(\mathcal{A})-P_{1}(\mathcal{A})$ where $\mathcal{A}=\{p_{0}>p_{1}\}$ is the Yatracos set with maximal discrepancy between $P_{0}$ and $P_{1}$. The witness function is $f^{*}_{(P_{0},P_{1})}(x)=\mathbb{I}(x\in\mathcal{A})-1/2$. An immediate advantage of targeting the TV projection comes from that $f^{*}$ is uniformly bounded. Given samples $\mathcal{D}$, consider the following test statistic which referred to as the split Scheffé statistic:
@@
Suppose $\rho(P_{0}\|P_{1})=\mathsf{W}_{1}({P_{0}},{P_{1}})$ is the 1-Wasserstein distance (or Kantorovich metric). The associated function class is =$\mathcal{F}=\{f:\left\|f\right\|_{L}\leq 1\}$ where $\left\|f\right\|_{L}:=\sup\{|f(x)-f(y)|/\|x-y\|:x\neq y\}$ is the Lipschitz semi-norm. Although the ideas are much more general, we limit our discussion to univariate distributions on a compact support, i.e., $\mathcal{X}=[0,b]$. In this case, the witness function is explicit and easy to describe [4].
@@
Suppose that $\mathcal{F}$ is a unit ball of the reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, with kernel $k(x,y)$, and RKHS norm $\|\cdot\|_{\mathcal{H}}$, i.e., =$\mathcal{F}=\{f:\left\|f\right\|_{\mathcal{H}}\leq 1\}$. Then the corresponding IPM is called the Maximum Mean Discrepancy [MMD, 26]. It was shown by Gretton et al. [26] that the analytic witness function $f^{*}_{(P,\widehat{P}_{1})}=\frac{\mu_{P}-\mu_{\widehat{P}_{1}}}{\left\|\mu_{P}-\mu_{\widehat{P}_{1}}\right\|_{\mathcal{H}}}$ where $\mu_{P}(\cdot):=\mathbb{E}_{P}[k(X,\cdot)]$ is the mean embedding of $P$.
The split statistic $\overline{T}_{n_{0}}(P,\widehat{P}_{1})$ in this case reduces to an average of the (negative) witness function $-P^{*}_{n_{0}}(f^{*}_{(P,\widehat{P}_{1})})$ if the kernel $k(\cdot,\cdot)$ is symmetric. In this case, the sign of the split statistic captures, in expectation, whether the population $P^{*}$ is closer to $P$ or $\widehat{P}_{1}$ based on mean embeddings.
@@ B.2 Example 3
[…]
@@ B.3 Example 4
[…]
@@
[…]
@@
[…]
@@ D.2 Proof of Proposition 4
[…]
@@ D.3 Proof of Proposition 6
[…]
@@
[…]
@@
$\widehat{P}_{1}=\begin{cases}\mbox{$\mathsf{Bern}$}(0),&\overline{X}_{n_{1}}<0.5-1/(2\sqrt{2})\approx 0.146\\ \mbox{$\mathsf{Bern}$}(1/2),&\text{Otherwise.}\\ \end{cases}$
@@ E.1 Proof of Theorem 5
Let the event $E$ be $\delta_{1}(\widehat{P}_{1},\widetilde{P})\leq(M/\nu)\eta_{n}$ which happens with probability at least $21-\kappa/$ 2. Then,
@@
Now we focus on bounding $\mathbb{E}_{P^{*}}[\sup_{P\in\mathcal{S}_{\epsilon}}|\widetilde{T}_{P}|]$ (which is greater than $\mathbb{E}_{P^{*}}[\sup_{P\in\mathcal{S}_{\epsilon}}\widetilde{T}_{P}]$). Let $\mathcal{F}_{T,\mathcal{P}}=\{T(\cdot;\widehat{P}_{1},P):P\in\mathcal{P}\}$. The symmetrization lemma [Lemma 2.3.1, 57] states that
@@ Appendix F Finite-sample valid confidence set for bounded test statistic
As in the case of Bernstein’s bound (31), Bentkus’s bound (33) requires prior knowledge of $P^{*}$ to compute the variance $S$. The empirical Bentkus’s bound [36] addresses this by taking the union bound on variance over-estimation and the Bentkus’s inequality. Following Kuchibhotla and Zheng [Lemma F.1, 36] define the over-estimator of $S$ as, for $\delta\in[0,1]$,
@@ Appendix G Crossfit ReDI set
$\mathbb{P}_{n_{0}}T(\cdot;P,\widehat{P}_{1})-\mathbb{P}_{P^{*}}T(\cdot;P,\widetilde{P})=$ $(\mathbb{P}_{n_{0}}-\mathbb{P}_{P^{*}})[T(\cdot;P,\widehat{P}_{1})-T(\cdot;P,\widetilde{P})]$
$+\mathbb{P}_{P^{*}}\left[T(\cdot;P,\widehat{P}_{1})-T(\cdot;P,\widetilde{P})\right]+(\mathbb{P}_{n_{0}}-\mathbb{P}_{P^{*}})T(\cdot;P,\widetilde{P}).$
@@
The $n(\overline{T}^{\times}(P)-\mathbb{P}_{P^{*}}T(\cdot;P,\widetilde{P}))$ has the following decomposition:
$n_{0}(\mathbb{P}_{n_{0}}-\mathbb{P}_{P^{*}})[T(\cdot;P,\widehat{P}_{1})-T(\cdot;P,\widetilde{P})]+n_{1}(\mathbb{P}_{n_{1}}-\mathbb{P}_{P^{*}})[T(\cdot;P,\widehat{P}_{0})-T(\cdot;P,\widetilde{P})]$
$+n_{0}\mathbb{P}_{P^{*}}[T(\cdot;P,\widehat{P}_{1})-T(\cdot;P,\widetilde{P})]+n_{1}\mathbb{P}_{P^{*}}[T(\cdot;P,\widehat{P}_{0})-T(\cdot;P,\widetilde{P})]$
$+n(\mathbb{P}_{n}-\mathbb{P}_{P^{*}})T(\cdot;P,\widetilde{P}).$
@@
where $\widehat{s}_{P}^{\times 2}=[\widehat{\mathbb{V}}_{P^{*}}(T(X;P,\widehat{P}_{1}))+\widehat{\mathbb{V}}_{P^{*}}(T(X;P,\widehat{P}_{0}))]/2$ is a consistent estimator of $\mathbb{V}_{P^{*}}(T(X;P,\widetilde{P}))$.
@@ I.1 Example 1
[…]
@@
[…]
@@ I.2 Example 2
$\widehat{p}_{n_{1}}=\begin{cases}\overline{X}_{n_{1}}\wedge 0.4,&\overline{X}_{n_{1}}<0.5\\ \overline{X}_{n_{1}}\vee 0.6,&\overline{X}_{n_{1}}>0.5\\ \end{cases}$
@@
Figure 13: $\rho(P^{*}\|\textrm{Poi}(\theta))$ across the varying dispersion ratio $\kappa$. Red + indicates $\rho(P^{*}\|\widetilde{P})$.
@@
As highlighted in Section 5.1, the split LRT confidence set $(\widehat{C}_{sLRT})$ may fail to cover the KL projection unlike the KL ReDI set $(\widehat{C}_{\mathsf{KL}\textsc{ReDI}})$ even with the same choice of $\widehat{\theta}_{1}$ and the same log split likelihood-ratio statistic. Figure 14 contrasts the performance of $\widehat{C}_{sLRT}$ and $\widehat{C}_{\mathsf{KL}\textsc{ReDI}}$ based on 1000 replicates of 200 simulated observations. In computing the confidence sets, the observations are equally split in half and we choose $\widehat{\theta}_{1}$ to be the sample mean (which is the MLE) of the first half samples. As the misspecification gets more severe (larger $\kappa$), the empirical coverage of KL projection parameter $(\widetilde{\theta})$ decreases for $\widehat{C}_{sLRT}$. When the dispersion ratio becomes larger than 3, $\widehat{C}_{sLRT}$ fails to achieve the nominal 95% coverage whereas $\widehat{C}_{\mathsf{KL}\textsc{ReDI}}$ maintains the validity regardless of how severe the misspecification is. Both the center and the right panel depict the size of the estimated confidence set varying over the dispersion ratio but from a different perspective. The former is based on the maximal excess KL divergence from the KL projection (which can be at most twice the KL-diameter of the set) whereas the latter is based on the $L_{2}$ distance over the parameter space. It is not surprising that compared to $\widehat{C}_{\mathsf{KL}\textsc{ReDI}}$, $\widehat{C}_{sLRT}$ is smaller in the $L_{2}$ sense and is closer to $\widetilde{P}$ in an excess divergence sense.
Figure 14: Inference on KL projection: (Left) Empirical coverage (Center) $\mathsf{KL}_{\mathscr{H}}^{P^{*}}\left(\widehat{C},\{\widetilde{P}\}\right)$ (Right) $L_{2}$ width of $\widehat{C}$ in the parameter space. Point estimates are median over 1000 replicates and the error bar indicates $\pm$ 1 standard error.
@@
Figure 16 compares the various TV ReDI set when the $P^{*}$ is a $32\%$ variance inflated negative binomial—Berry-Esseen ($\widehat{C}_{\mathsf{TV}\textsc{ReDI}}$), Hoeffding bound ($\widehat{C}_{\mathsf{TV}\textsc{HF}}$), empirical Bernstein bound [$\widehat{C}_{\mathsf{TV}\textsc{EBS}}$; 59], and empirical Bentkus bound [$\widehat{C}_{\mathsf{TV}\textsc{EBK}}$; 9, 36]. See Section F for explicit forms of each confidence set. In all cases, we choose the same minimum TV distance estimator $\widehat{\theta}_{1}$. The KL ReDI set dominates all finite sample valid confidence sets considered in this section, despite its validity relying on asymptotics. The finite sample valid ReDI sets are too conservative (and yield a meaningless set $\widehat{C}=\mathcal{P}$) when only a few observations are available ($n\leq 50$). Although our paper does not primarily focus on obtaining the tightest finite-sample valid confidence set, leveraging the variance $\mathbb{V}_{P^{*}}(X)$ can often be beneficial when constructing the confidence set. In this example, $\widehat{C}_{\mathsf{TV}\textsc{EBS}}$ and $\widehat{C}_{\mathsf{TV}\textsc{EBK}}$ outperform $\widehat{C}_{\mathsf{TV}\textsc{HF}}$ since the Bernstein and Bentkus bounds are more sensitive to the variance.
@@
We adopt a heuristic search method for finding a confidence set in multivariate parameter space. For brevity, we explain the procedure in 2 dimensions, but the procedure can be straightforwardly extended to higher dimensions. From the observation that when $\widehat{\theta}_{1}$ is close to $\widetilde{\theta}$, i.e., $\rho(P^{*}\|P_{\widehat{\theta}_{1}})\leq\nu\rho(P^{*}\|\widetilde{P})$ as seen in the proof of Theorem 2, $\overline{T}_{n_{0}}(\widehat{\theta}_{1})=0$ for split statistics that satisfies Assumption 1 and 2. Therefore, we construct a star-convex confidence set that always includes $\widehat{\theta}_{1}$. We construct the rays originate from $\widehat{\theta}_{1}$, i.e., $R_{\omega}=\{\theta\in\Theta:r_{\omega}^{\top}(\theta-\widehat{\theta}_{1})=0,r\geq 0\}$ where $r_{\omega}=(r\sin\omega,-r\cos\omega)$ for angle $\omega\in[-\pi,\pi]$. For each $\omega$, we find a root of an evidence function $\operatorname*{ev}(\theta)=\overline{T}_{n_{0}}(\theta)-t_{\alpha}(\theta)$ using Brent’s method [13] on $R_{\omega}$ constrained with radius $r$ varying from 0 (corresponding $\theta=\widehat{\theta}_{1}$) to some $r_{0}>0$ such that the corresponding $\widetilde{\theta}$ satisfies $\operatorname*{ev}(\widetilde{\theta})>0$.
@@
Figure 22 shows the average maximum width of the confidence set. Here we denote maximum width since the resulting confidence set may be the union of 00{0} and some interval not including 00. The width of confidence set the largest for DPReDI with $\beta=1$, followed by DPReDI with $\beta=0.1$, KLReDI, sLRT, LRT1 and LRT2. This ordering underscores the conservatism of (robust) universal inference method as well as of choosing robust target of inference—The higher the $\beta$ is, the more robust the DP projection becomes—when the model is well specified. In all (robust) universal inference methods, crossfit reduced the size of the confidence set without sacrificing the validity. Particularly, the width of the crossfit KLReDI is comparable to vanilla sLRT. This result precisely matches with the power comparison under Gaussian location model in the Supplementary Material S8.1.

### v4

@@ 10 Discussion
The papers [59, 24] study many variants of universal inference (including constructing confidence sequences instead of confidence sets, to combining multiple sample-splits) and investigating these variants in the context of the robust universal inference framework of this paper would be interesting. One can apply $\mathsf{KL}\textsc{ReDI}$ to estimate the KL projection conditional distribution under model misspecification. See Supplementary Material M for details. Extending these results to other projections in the regression context would be another interesting future work.
@@ Appendix A Examples of IPM-ReDI set
Suppose $\rho(P_{0}\|P_{1})=\mathsf{TV}(P_{0},P_{1})$ where $\mathsf{TV}$ is the total variation distance. This is an IPM over the function class $\mathcal{F}=\{f:\left\|f\right\|_{\infty}\leq 1\}$. An equivalent definition is $\mathsf{TV}(P_{0},P_{1})=\sup_{A}|P_{0}(A)-P_{1}(A)|=P_{0}(\mathcal{A})-P_{1}(\mathcal{A})$ where $\mathcal{A}=\{p_{0}>p_{1}\}$ is the Yatracos set with maximal discrepancy between $P_{0}$ and $P_{1}$. The witness function is $f^{*}_{(P_{0},P_{1})}(x)=\mathbb{I}(x\in\mathcal{A})-1/2$. An immediate advantage of targeting the TV projection comes from that $f^{*}$ is uniformly bounded. Given samples $\mathcal{D}$, consider the following test statistic which referred to as the split Scheffé statistic:
@@
Suppose $\rho(P_{0}\|P_{1})=\mathsf{W}_{1}({P_{0}},{P_{1}})$ is the 1-Wasserstein distance (or Kantorovich metric). The associated function class is $\mathcal{F}=\{f:\left\|f\right\|_{L}\leq 1\}$ where $\left\|f\right\|_{L}:=\sup\{|f(x)-f(y)|/\|x-y\|:x\neq y\}$ is the Lipschitz semi-norm. Although the ideas are much more general, we limit our discussion to univariate distributions on a compact support, i.e., $\mathcal{X}=[0,b]$. In this case, the witness function is explicit and easy to describe [4].
@@
Suppose that $\mathcal{F}$ is a unit ball of the reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, with kernel $k(x,y)$, and RKHS norm $\|\cdot\|_{\mathcal{H}}$, i.e., $\mathcal{F}=\{f:\left\|f\right\|_{\mathcal{H}}\leq 1\}$. Then the corresponding IPM is called the Maximum Mean Discrepancy [MMD, 27]. It was shown by Gretton et al. [27] that the analytic witness function $f^{*}_{(P,\widehat{P}_{1})}=\frac{\mu_{P}-\mu_{\widehat{P}_{1}}}{\left\|\mu_{P}-\mu_{\widehat{P}_{1}}\right\|_{\mathcal{H}}}$ where $\mu_{P}(\cdot):=\mathbb{E}_{P}[k(X,\cdot)]$ is the mean embedding of $P$.
The split statistic $\overline{T}_{n_{0}}(P,\widehat{P}_{1})$ in this case reduces to an average of the (negative) witness function $P^{*}_{n_{0}}(f^{*}_{(P,\widehat{P}_{1})})$ if the kernel $k(\cdot,\cdot)$ is symmetric. In this case, the sign of the split statistic captures, in expectation, whether the population $P^{*}$ is closer to $P$ or $\widehat{P}_{1}$ based on mean embeddings.
@@ B.2 Example 3
$\frac{\partial}{\partial\theta}\mathsf{KL}(P^{*}\|P_{\theta})=\int\frac{\partial}{\partial\theta}\log\left(\frac{p^{*}(x)}{p_{\theta}(x)}\right)\,{\mathrm{d}}P^{*}(x)\propto\int(x-\theta)\,{\mathrm{d}}P^{*}=\epsilon_{n}a_{n}-\theta\overset{\text{set}}{=}0$
@@ B.3 Example 4
The KL projection $\widetilde{P}$ is $\mbox{$\mathsf{Bern}$}\left(3/4\right)$. For simplicity we suppose that $n$ is even, and that $\mathcal{D}_{0}$ consists of the first $n/2$ samples and $\mathcal{D}_{1}$ consists of the remaining samples. For a constant $\beta>0$, let us consider the events $E_{0},E_{1}$ defined as,
@@
When events $E_{0}$ and $E_{1}$ hold we can see that the universal set $C_{\alpha}(X_{1},\ldots,X_{n})$ fails to cover $\widetilde{P}$. In more detail, on the event $E_{1}$ the MLE, $\widehat{P}_{1}$ is $\mbox{$\mathsf{Bern}$}(1/4)$ and thus,
@@
we can reason instead about the probability of the events $E_{0}$ and $E_{1}$ when drawing samples from $\mbox{$\mathsf{Bern}$}(1/2)$, and account for the difference using the Total Variation. Combining this fact with the standard Berry-Esseen bound applied to Bernoulli sums, together with some simple algebraic manipulations, we obtain that for some universal constant $C>0$,
@@ D.2 Proof of Proposition 4
Since $\epsilon_{n}<(1-\alpha)/n$, the projection will be $\mbox{$\mathsf{Bern}$}(0)$ for any $\beta>0$, provided that $n\geq(1-\alpha)(1+\beta)/(1-(1/2)^{\beta})$.
@@ D.3 Proof of Proposition 6
Note that Hellinger projection is $\mbox{$\mathsf{Bern}$}(0)$ for $n>6$ (as long as $\epsilon_{n}<0.146$) since
@@
Similarly, $\nu$-approximate Hellinger projection is $\mbox{$\mathsf{Bern}$}(0)$ if $\epsilon_{n}<0.051$ or $\mathcal{P}$ otherwise. Hereafter we only consider $n>20$ where $\widetilde{P}_{\nu}=\widetilde{P}$.
@@
[…]
@@ E.1 Proof of Theorem 5
Let the event $E$ be $\delta_{1}(\widehat{P}_{1},\widetilde{P})\leq(M/\nu)\eta_{n}$ which happens with probability at least $1-\kappa/2$. Then,
@@
Now we focus on bounding $\mathbb{E}_{P^{*}}[\sup_{P\in\mathcal{S}_{\epsilon}}|\widetilde{T}_{P}|]$ (which is greater than $\mathbb{E}_{P^{*}}[\sup_{P\in\mathcal{S}_{\epsilon}}\widetilde{T}_{P}]$). Let $\mathcal{F}_{T,\mathcal{P}}=\{T(\cdot;\widehat{P}_{1},P):P\in\mathcal{P}\}$. The symmetrization lemma [Lemma 2.3.1, 58] states that
@@ Appendix F Finite-sample valid confidence set for bounded test statistic
As in the case of Bernstein’s bound (31), Bentkus’s bound (33) requires prior knowledge of $P^{*}$ to compute the variance $S$. The empirical Bentkus’s bound [37] addresses this by taking the union bound on variance over-estimation and the Bentkus’s inequality. Following Kuchibhotla and Zheng [Lemma F.1, 37] define the over-estimator of $S$ as, for $\delta\in[0,1]$,
@@ Appendix G Crossfit ReDI set
[…]
@@
The $)n(\overline{T}^{\times}(P)-\mathbb{P}_{P^{*}}T(\cdot;P,\widetilde{P}$)) has the following decomposition:
[…]
@@
where $\widehat{s}_{P}^{\times 2}=[\widehat{\mathbb{V}}_{P^{*}}(T(X;P,\widehat{P}_{1}))+\widehat{\mathbb{V}}_{P^{*}}(T(X;P,\widehat{P}_{0}))]/2$ is a consistent estimator of $)\mathbb{V}_{P^{*}}(T(X;P,\widetilde{P}$)).
@@ I.1 Example 1
DP Projection Recall that DP projection is $\mbox{$\mathsf{Bern}$}(0)$ for small enough $\epsilon_{n}$ Following the proof of Proposition 4. We only consider DP parameter $\beta<2$ here.
@@
Hellinger Projection Given $\epsilon\leq 0.04$, (approximate) Hellinger projection $\widetilde{P}_{\sf H}$ is $\mbox{$\mathsf{Bern}$}(0)$.
@@ I.2 Example 2
[…]
@@
[…]
@@
[…]
@@
Figure 16 compares the various TV ReDI set when the $P^{*}$ is a $32\%$ variance inflated negative binomial—Berry-Esseen ($\widehat{C}_{\mathsf{TV}\textsc{ReDI}}$), Hoeffding bound ($\widehat{C}_{\mathsf{TV}\textsc{HF}}$), empirical Bernstein bound $\widehat{C}_{\mathsf{TV}\textsc{EBS}}$; 60], and empirical Bentkus bound $\widehat{C}_{\mathsf{TV}\textsc{EBK}}$; 9, 37]. See Section F for explicit forms of each confidence set. In all cases, we choose the same minimum TV distance estimator $\widehat{\theta}_{1}$. The KL ReDI set dominates all finite sample valid confidence sets considered in this section, despite its validity relying on asymptotics. The finite sample valid ReDI sets are too conservative (and yield a meaningless set $\widehat{C}=\mathcal{P}$) when only a few observations are available ($n\leq 50$). Although our paper does not primarily focus on obtaining the tightest finite-sample valid confidence set, leveraging the variance $\mathbb{V}_{P^{*}}(X)$ can often be beneficial when constructing the confidence set. In this example, $\widehat{C}_{\mathsf{TV}\textsc{EBS}}$ and $\widehat{C}_{\mathsf{TV}\textsc{EBK}}$ outperform $\widehat{C}_{\mathsf{TV}\textsc{HF}}$ since the Bernstein and Bentkus bounds are more sensitive to the variance.
@@
[…]
@@
Figure 22 shows the average maximum width of the confidence set. Here we denote maximum width since the resulting confidence set may be the union of 0{0} and some interval not including 0. The width of confidence set the largest for DPReDI with $\beta=1$, followed by DPReDI with $\beta=0.1$, KLReDI, sLRT, LRT1 and LRT2. This ordering underscores the conservatism of (robust) universal inference method as well as of choosing robust target of inference—The higher the $\beta$ is, the more robust the DP projection becomes—when the model is well specified. In all (robust) universal inference methods, crossfit reduced the size of the confidence set without sacrificing the validity. Particularly, the width of the crossfit KLReDI is comparable to vanilla sLRT. This result precisely matches with the power comparison under Gaussian location model in the Supplementary Material S8.1.
@@
Appendix M KLReDI for Conditional Estimation
Suppose we observe i.i.d. paired samples $(X_{1},Y_{1}),\dots(X_{n},Y_{n})\in\mathcal{X}\times\mathcal{Y}$ drawn from a joint distribution $P^{*}_{X,Y}\in\mathcal{Q}$ where $P^{*}_{Y|X}=P^{*}_{Y|X}\otimes P^{*}_{X}$, $P^{*}_{Y|X}\in\mathcal{Q}_{Y|X}$ is the true conditional response distribution, and $P^{*}_{X}\in\mathcal{Q}_{X}$ is a true marginal regressor distribution following the notation from Buja et al. [14]. We assume that $\mathcal{Q}_{Y|X}$ is a class of regular conditional probability distributions. Suppose the working model $\mathcal{P}_{Y|X}$ is a collection of (regular) conditional response distributions parametrized by $\theta$, i.e., $\mathcal{P}_{Y|X}=\{P_{Y|X;\theta}:\theta\in\Theta\}$. Let the marginal regressor distribution $P_{X}\in\mathcal{M}(\mathcal{X})$ be any probability distribution on $\mathcal{X}$. Then the collection of the joint probability distribution from $\mathcal{P}_{Y|X;\Theta}$ and $\mathcal{M}(\mathcal{X})$ is given by $\mathcal{P}_{\Theta}=\{P_{Y|X;\theta}\otimes P_{X}:\theta\in\Theta,P_{X}\in\mathcal{M}(\mathcal{X})\}$.
Suppose our target of inference is the KL projection parameter defined as
$\widetilde{\theta}=\operatorname*{arg\,inf}_{\theta\in\Theta,P_{X}\in\mathcal{M}(\mathcal{X})}\mathsf{KL}\left(P^{*}_{X,Y}\|P_{Y|X;\theta}\otimes P_{X}\right)=\operatorname*{arg\,inf}_{\theta\in\Theta}\mathsf{KL}\left(P^{*}_{Y|X}\|P_{Y|X;\theta}\right),$
where the last equality is due to the fact that the infimum over $\mathcal{M}(\mathcal{X})$ is attained at $P_{X}=P^{*}_{X}$ irrespective of the choice $\theta$:
$\inf_{P_{X}\in\mathcal{M}(\mathcal{X})}\mathsf{KL}(P^{*}_{X,Y}\|P_{Y|X}\otimes P_{X})$ $=\inf_{P_{X}}\int\log\frac{P^{*}_{X,Y}}{P_{Y|X}\otimes P_{X}}\,{\mathrm{d}}P^{*}_{X,Y}$
$=\int\log\frac{P^{*}_{Y|X}}{P_{Y|X}}\,{\mathrm{d}}P^{*}_{Y|X}+\inf_{P_{X}}\int\log\frac{P^{*}_{X}}{P_{X}}\,{\mathrm{d}}P^{*}_{X}$
$=\mathsf{KL}(P^{*}_{Y|X}\|P_{Y|X}).$
To design a $\mathsf{KL}\textsc{ReDI}$ set of $\widetilde{\theta}$, we follow steps analogous to those in Section 5.1 using the natural plug-in estimator of the difference in KL divergence. In particular, we can write the difference in the KL divergence as follows:
$\mathsf{KL}(P^{*}_{X,Y}\|P_{Y|X;\theta}\otimes P_{X})-\mathsf{KL}(P^{*}_{X,Y}\|P_{Y|X;\widehat{\theta}_{1}}\otimes P_{X})$ $=\int\log\frac{P_{Y|X;\widehat{\theta}_{1}}}{P_{Y|X;\theta}}\,{\mathrm{d}}P^{*}_{X,Y}.$
Therefore, the natural (split) plug-in estimator of the difference in KL divergence is
$\overline{T}_{n_{0}}=\frac{1}{n_{0}}\sum_{i\in\mathcal{D}_{0}}T_{i}(\theta,\widehat{\theta}_{1}),\qquad T_{i}(\theta,\widehat{\theta}_{1})\equiv T\left((X_{i},Y_{i});\theta,\widehat{\theta}_{1}\right)=\log\frac{p_{Y|X_{i};\widehat{\theta}_{1}}(Y_{i})}{p_{Y|X_{i};\theta}(Y_{i})},$
where $\widehat{\theta}_{1}$ is a pilot estimate constructed from the split sample $\mathcal{D}_{1}$. We can now construct the $\mathsf{KL}\textsc{ReDI}$ set $\widehat{C}_{\mathsf{KL}\textsc{ReDI},n}$ with $\delta$-corrupted samples analogous to (17).